OpenStudy (anonymous):

write the polar equation in rectangular form r=-12 cos theta

myininaya (myininaya):

hint multiply r on both sides

OpenStudy (anonymous):

so -12 times r?

myininaya (myininaya):

\[r^2=-12 r \cos(\theta)\]

myininaya (myininaya):

I bet you know what rcos(theta) equals and r^2 equals in terms of x and y.

OpenStudy (anonymous):

im lost sorry

myininaya (myininaya):

\[x^2+y^2=r^2 \\ rcos(\theta)=x \\ rsin(\theta)=y\]

myininaya (myininaya):

\[\tan(\theta)=\frac{y}{x}\]

myininaya (myininaya):

you only need two of the equations I listed

myininaya (myininaya):

for this particular equation

OpenStudy (anonymous):

so this is the answer?

myininaya (myininaya):

what is?

OpenStudy (anonymous):

tan(0)=x/y

myininaya (myininaya):

how did you get

myininaya (myininaya):

\[r^2=-12 r \cos(\theta)\] do you see the r^2 and the rcos(theta)?

OpenStudy (anonymous):

tan(theta)

myininaya (myininaya):

I gave you two equations above that told you what those were in terms of x and y

myininaya (myininaya):

do you want to try again to apply those equations I gave to your equation

myininaya (myininaya):

\[r=-12 \cos(\theta) \\ r \cdot r =r \cdot (-12 \cos(\theta)) \\ r^2=-12 r \cos(\theta) \\ \text{ replace } r^2 \text{ with } x^2+y^2 \\ \text{ replace } r \cos(\theta) \text{ with } x \\ \text{ this is \because } r^2=x^2+y^2 \text{ and } r \cos(\theta)=x \]